OpenStudy (anonymous):

Simplify the expression. 6*2^-3

OpenStudy (anonymous):

\[X ^{-b}=\frac{ 1 }{ x ^{b} }\]

OpenStudy (anonymous):

whats next/

OpenStudy (anonymous):

\[6*2^{-3}=\frac{ 6 }{ 2^{3} }\]

OpenStudy (anonymous):

and 2^3=2*2*2=8 the final ans 6/8=3/4

OpenStudy (anonymous):

ohh! so thats the formula i do everytime for this kind of problem? thank you for the steps